OpenStudy (pinklion23):

slope help 2 questions pls

OpenStudy (pinklion23):

find the slope y-2=-3(x+1)

OpenStudy (pinklion23):

@escamer

OpenStudy (escamer):

hmmm, i learned this recently so i might be able to do it

OpenStudy (pinklion23):

@imqwerty

OpenStudy (escamer):

@RCCB

pooja195 (pooja195):

find the slope y-2=-3(x+1) Start by isolating y

pooja195 (pooja195):

-3(x+1) Distribute

OpenStudy (pinklion23):

y=-3(x+1)+2?

pooja195 (pooja195):

-3 x X =? -3 x 1 =?

OpenStudy (pinklion23):

-3*x=-3x -3*1= -3

pooja195 (pooja195):

ok y=mx+b y-2=-3x-3 Without doing anything else we already know the slope m=slope

OpenStudy (pinklion23):

?

pooja195 (pooja195):

y=-3x-1 y=mx+b <---slope intercept form m=slope b=-y intercept Figure out the slope (its the number next to the x)

OpenStudy (pinklion23):

-1?

OpenStudy (pinklion23):

er -3

OpenStudy (pinklion23):

*brain fart*

OpenStudy (pinklion23):

imwerty? what do you think :P

pooja195 (pooja195):

-3 is correct

OpenStudy (pinklion23):

ok sorry and thx

OpenStudy (pinklion23):

i have one more

OpenStudy (pinklion23):

\[y+3=-\frac{ 3 }{ 4 }(x-4)\]

OpenStudy (pinklion23):

brb

pooja195 (pooja195):

Just combine -3/4 with x and thats your slope.

OpenStudy (pinklion23):

ok im back so the slope is -3/4x?

pooja195 (pooja195):

Yes its just -3/4

pooja195 (pooja195):

Dont add the x

OpenStudy (pinklion23):

ok thank you so much :)